Tohban Report 2014-05-28

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 21 May 2014
End Date: 28 May 2014
Tohban: Juan Carlos Martinez Oliveros (Vice Tohban: Martin Fivian)
Tohban email:
Next Tohban: Nicole Duncan
List all reports


Solar Activity

Solar activity this week was low. There are currently 6 active regions on the disk. Farside monitoring shows an active region that may rotate into the disk in the 5-6 days.

How many GOES flares occurred?

 Flares above B, C, M, X class were      9    27     1     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      2    16     0     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      2     6     0     0

There were RHESSI flares/GOES flares 63 / 37 over the time range 21-May-14 28-May-14

Memory Management

The SSR was not emptying by the end of the daily pass set, reaching a maximum of 53% and a minimum of about 12%

Spacecraft Status

The cold plate 1 and 2 are at 118.4 K and 116.9 K respectively.

Data Gaps

No data gaps this week.

Detector issues

On may 22, detector 6 HV was decreased by 20 counts and detector 9 front fast threshold from 0x40 to 0x50

  2014-142-21:31:54 /IDPULOAD VALUE=0x50         ; was 0x40
  2014-142-21:39:28 /IHVDAC DETECTOR=6, VOLTAGE=88 ; was 108

On May 26 also the following thresholds were changed:

  14-147-16:56:13 /IDPULOAD VALUE=0x30
  14-147-16:56:34 /IDPULOAD VALUE=0x30
  14-147-16:57:00 /IDPULOAD VALUE=0x30
  14-147-16:57:26 /IDPULOAD VALUE=0x3A

On May 26 detector 3 GV was decreased by 20 counts

  14-147-20:17:56 /IHVDAC DETECTOR=3, VOLTAGE=186

On 2014 May 28 the G9 experiment was made. The information about the experiment is here [1].

Other notes

The thin shutter came in 2014-05-25 23:50 and stayed in. The issue was solved on Monday. After a series of front detector threshold changes.

Spacecraft Management

A spin up to 15 RPM was made on 2014-141-20:22 UTC and returned to normal mode at 21:32 UTC.

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) +/- 4 minutes
Night time data (rears) +/- 4 minutes
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Active
Detector problems? Many adjustments to Det settings; see above.
Personal tools